# PlanQA: A Benchmark for Spatial Reasoning in LLMs using Structured Representations

### v1

@@ Appendix F Failure Cases
F.1 Max Box
As illustrated in Figure 10, this example shows GPT-4.1 producing a suboptimal result in a spatial reasoning task. Instead of identifying the largest free region in the kitchen layout, the model placed a bounding box that extended along a single axis, starting from one edge, without fully exploring the layout. It selected an area from x=0 to 2.1 and y=1.1 to 1.8, yielding 1.47 m² of free space.
A larger available region, shown in Figure 10, spans from (0.7, 0.7) to (2.1, 1.8), and offers 1.54 m². The model did not identify this larger region, which indicates that its search strategy was likely not exhaustive, even though it had access to the full structured input. The reasoning trace in Figure 11 confirms that it evaluated a set of candidate regions but ultimately settled on a suboptimal one. The criteria for candidate selection remain unclear and may vary across queries.
Figure 9: GPT-4.1 prediction
Figure 10: Ground truth
LLM Reasoning Summary Let’s analyze the problem step by step. First, I’ll look for large free areas without obstacles. Option 1: The region to the right of the chairs and table, from x=1.3 to 2.1, and from y=0 to 1.8. This gives an estimated area of 1.44 m². Option 2: The space above the chairs and below the wall units, from x=0 to 2.1 and y=1.1 to 1.8. This gives an area of 1.47 m². Option 3: A small gap near the bottom right corner. It looks too narrow and likely blocked by furniture. Option 4: The space in front of chair_2, but it appears too small to be useful. Based on these, Option 2 seems best. I’ll choose the rectangle [0.0, 1.1, 2.1, 1.8], which gives an area of 1.47 m².
Figure 11: Summarization of chain-of-thought style output from GPT-4.1. The model imitates reasoning by listing and evaluating candidate regions, then selects a suboptimal solution.
F.2 Shortest Path
In this task, the model must find a path from the center of a starting object to the center of a destination object. For simplicity, the start and goal objects are not treated as obstacles.
Our pathfinding system constructs a visibility graph by connecting all mutually visible corner points in the environment. Dijkstra’s algorithm is then used to compute the true shortest path over this graph, ensuring optimality under Euclidean distance. This method avoids the limitations of grid-based approximations and produces precise, flexible paths around obstacles. While multiple valid solutions may exist depending on object layout and clearance constraints, the chosen path always reflects the shortest feasible route.
When interpreting the task, models often fail to make effective use of available free space. Instead of navigating around obstacles or adapting to the environment, they frequently extend a straight line in one direction until it is blocked. This leads to incomplete or inefficient trajectories that do not reflect the geometry of the room or required clearance. In some cases, paths cut too closely around objects, likely due to the model implicitly minimizing Euclidean distance without accounting for spatial constraints.
Figure 13 shows a path that reaches the target but fails the Fréchet distant constraint of 0.6 m, making it non-optimal. Figure 13 depicts a path that violates clearance constraints and does not represent a valid trajectory to the target.
Figure 12: Comparison between the optimal path (green) and the GPT-4.1-generated trajectory (cyan), which, despite reaching the target, exceeds the Fréchet distance threshold of 0.6 m.
Figure 13: Depiction of the optimal path (green) alongside the GPT-4.1-generated trajectory (red), which fails to satisfy clearance requirements and results in an infeasible route.
Appendix G VLM Experiments
To further probe model performance on particularly challenging tasks, we conducted targeted evaluations of two vision-language models (VLMs). We focused on the most difficult question categories–Free Space and Max Box–within the most spatially complex environments: living rooms and bedrooms. The models tested in this focused benchmark include Gemini 2.5 Flash Preview 04-17 Gemini Team, Google (2025) and GPT-4.1. Results are summarized in Table 12, which compares each model’s output under two input conditions: JSON only and JSON combined with a floorplan image. An example of Gemini’s step-by-step reasoning under the JSON+Image condition is shown in Figure 14.
Table 12: Comparison of spatial understanding between Gemini 2.5 Flash and GPT-4.1 for Living Room (L) and Bedroom (B), under different conditions.
Type Category [Uncaptioned image] Gemini 2.5 Flash [Uncaptioned image] GPT-4.1
JSON Only JSON + Image JSON Only JSON + Image
L Free Space 6.4 12.2 5.0 9.8
Max Box 17.0 22.0 6.5 4.3
B Free Space 12.8 15.2 2.8 10.5
Max Box 15.7 21.0 7.2 7.2
[…]
Figure 14: Summarized step-by-step reasoning by Gemini to compute maximum free space from a JSON floorplan and image. The model merges nested objects and analyzes rectangular gaps to estimate usable area.
While our results show that both models benefit from visual input in some settings, the extent of improvement varies by task. For the Free Space category, both Gemini 2.5 Flash and GPT-4.1 show substantial gains when provided with floorplan images. In contrast, for Max Box, only Gemini shows a notable improvement, while GPT-4.1 performs similarly or worse with visual input. This variation suggests that visual grounding supports some spatial tasks more effectively than others.
The stronger visual performance of Gemini 2.5 Flash may stem from architectural or training differences, possibly reflecting design choices aimed at robotics or spatial reasoning tasks.
One challenge lies in the nature of floorplan inputs. Unlike natural images, floorplans vary widely in style, abstraction, and labeling. This inconsistency makes it harder for models to extract spatial meaning reliably.
Smaller models show little benefit from visual input, likely due to limited capacity and lack of pretraining for structured spatial tasks.
In summary, floorplan understanding requires more than object detection. It demands precise spatial reasoning and integration with structured formats like JSON. Without task-specific tuning or standardized visual formats, even advanced models remain limited in this domain.
@@ Appendix H Prompts
This section presents selected prompt templates used in our benchmark. We include a summarized version of the bedroom layout generation prompt, as well as representative examples of question prompts targeting specific spatial reasoning skills. These prompts reflect the structured task design outlined in Table 11. Additional prompts–including full versions, formatting logic, and implementation details–are provided in the supplementary code and materials for reproducibility.
Prompt: Max Box You’ll be given a layout of a {room_type} in JSON format. Your task is to calculate the area (in m²) of the largest axis-aligned rectangle that can fit within the room. The rectangle must not overlap with any other objects, but it may fully occupy free space, including areas above rugs. You can ignore passageways, door clearance, or accessibility constraints. Room layout: {room_json} Reason step by step to calculate the area. Respond only with a float value in square meters using the format: *Final answer*: area
@@
Prompt: Missing Object Placement You’ll be given a layout of a {room_type} in JSON format. Your task is to suggest a suitable placement for the missing object {object_name} by providing its bounding box coordinates. The placement must avoid overlaps, respect obstacles, and maintain accessibility. Room layout: {room_json} Reason step by step to identify the placement. Respond only with bounding box coordinates (float values) using the format: *Final answer*: [x_min, y_min, x_max, y_max]
Figure 16: Prompt for placing a missing object in a room layout using Chain-of-Thought reasoning. The model must return valid coordinates that satisfy geometric constraints and maintain functional space.
[…]

### v2

@@ Appendix G Case Studies by Question Type
G.1 Pair Distance
Task Definition
In this task, the model is asked: “Calculate the euclidean distance between the centroids of the obj_1 and the obj_2.” In the visualization in Figure 8, these two polygons correspond to the sink and the shower; the goal is to compute their centroid-to-centroid distance.
Figure 7: Pair Distance (bathroom): the red line indicates the centroid-to-centroid segment.
Figure 8: Repositioning (bedroom): red fill indicates the initial pose of bin_2 and green fill indicates the final pose.
Ground-Truth Computation
To establish the correct answer, we first compute the centroid $(x,y)$ of each polygon. This is done using the shoelace formula, which calculates the centroid based on the polygon’s vertices. Once both centroids are obtained, the Euclidean distance between them is computed. A predicted answer is considered correct if it falls within a tolerance of $2\%$ of the ground-truth distance.
Main Issue
Models often fail on the HSSD dataset because they compute the centroid incorrectly. In earlier experiments, some models used the center of mass instead of the centroid; therefore, the word centroid is now explicitly stated in the prompt. Almost all wrong answers come from calculation mistakes in the centroid formula (areas, sums, divisions), not from the distance step. This error does not depend on polygon complexity (number of vertices).
G.2 Repositioning
Task Definition
We pose the question: “Calculate how far the object can be moved in the direction before it touches another object or the wall.” In the visualization in Figure 8, the object of interest is bin_2 and the direction is leftward; the goal is to compute its maximum leftward translation before contacting a bathtub.
Ground-Truth Computation
Simulate a leftward, axis-aligned slide of bin_2. Advance until the next step would overlap a wall or another object; take the last non-overlapping pose. Measure the travel distance from the initial position to that pose. Accept model answers within $2\%$ tolerance.
Main Issue
Narrow gaps and obstacles with arbitrary orientations make clearance difficult to estimate. Models often fall back to axis-aligned bounding boxes (discarding shape orientation) or omit the obstacle-union step, which leads to systematic over- or underestimation of feasible travel distances in any direction (left, right, up, or down).
G.3 Free Space
Task Definition
We consider the following question: “Calculate the total non-occupied floor area in the room?” In the visualization in Figure 10, the mint-highlighted area represents the space that remains free of objects within the game room; the goal is to compute its total area.
Figure 9: Unoccupied Floor Area (game room): the mint fill indicates the unoccupied (free-space) region within the game room.
Figure 10: View Angle (kitchen): smallest absolute angle between the vector from the centroid of the chair_2 to the centroid of the window_1 and global north $(0,1)$.
Ground-Truth Computation
We compute the free floor area using geometric operations provided by Shapely (Gillies et al., 2007). All object polygons within the room are merged using unary_union to correctly handle overlaps. The occupied area is then obtained from this union, and the free area is computed as
$A_{\text{free}}=A_{\text{room}}-A_{\text{union(objects)}}.$
A model prediction is considered correct if it falls within $5\%$ of the ground-truth free area.
Main Issue
Accurate free-space estimation hinges on correctly handling overlapping obstacles. A common failure mode is to subtract each object’s area independently, rather than forming their geometric union, which double-counts overlaps and systematically underestimates available area. For instance, when we evaluate on HSSD layouts using GPT-OSS-120B, cumulative accuracy declines as object count and overlap increase (see Figure 11), consistent with this union-omission error.
Figure 11: Cumulative accuracy versus layout complexity for HSSD using GPT-OSS-120B on the Free-Space task. Accuracy declines sharply as object count and overlap increase, reflecting the model’s difficulty in handling overlapping geometries.
G.4 View Angle
Task Definition
The prompt is: “Compute the smallest absolute angle between the vector from the centroid of obj_1 to the centroid of obj_2 and the global north vector $(0,1)$; report $\theta$ in degrees.” In the visualization in Figure 10 (kitchen), obj_1 is the chair_2 and obj_2 is the window_1; the goal is to return $\theta$, judged correct within a $2\%$ tolerance.
Ground-Truth Computation
(1) Compute centroids $\mathbf{c}_{s}$ (for sofa) and $\mathbf{c}_{v}$ (for TV) using the polygon shoelace formula.
(2) Form the displacement vector $\mathbf{d}=\mathbf{c}_{v}-\mathbf{c}_{s}$ and its unit vector $\hat{\mathbf{d}}=\dfrac{\mathbf{d}}{\lVert\mathbf{d}\rVert}$.
(3) Let the global north vector be $\mathbf{n}=(0,1)$ (already unit length). Compute the cosine via $\cos\theta=\mathrm{clip}\!\big(\hat{\mathbf{d}}\cdot\mathbf{n},\,-1,\,1\big)$.
(4) Convert to degrees and take the smallest absolute angle: $\theta=\arccos(\cos\theta)\cdot\dfrac{180}{\pi}\ \in[0^{\circ},180^{\circ}]$.
A prediction is correct if it is within $2\%$ of the ground-truth angle $\theta$.
Main Issue
On HSSD layouts, most errors come from centroid calculation mistakes (areas/sums/divisions in the shoelace step), not from the dot product. To avoid ambiguity, the prompt explicitly says centroid. On synthetic layouts (4-point, axis-aligned boxes), centroids are trivial, and this issue does not appear.
G.5 Placement
Task Definition
We consider the following question: “Check if a given object can be placed in the room without overlapping walls or other objects?” In the visualization in Figure 13, the object is a $2.5\,\mathrm{m}\times 1.0\,\mathrm{m}$ antique storage chest and the room is the living room; the goal is to determine whether a collision-free placement is possible. This task evaluates collision detection, spatial constraints, and free-space reasoning.
Figure 12: Placement (living room): determine whether an antique storage chest can be placed without overlapping walls or existing objects (collision-free feasibility).
Figure 13: Shortest Path (living room): ground-truth shortest walkable path with $15\,\mathrm{cm}$ clearance is shown in purple; the model’s predicted path in red intersects armchair_1, illustrating a failure case for GPT-OSS-120B due to incorrect obstacle/clearance handling.
Ground-Truth Computation
Form the room polygon and the union of all existing object polygons. Allow arbitrary rotation (non–axis-aligned) for the $2\times 1.5$ m rectangle. Search over poses: for each orientation $\theta$, test placements where the rotated rectangle is strictly inside the room polygon and has no intersection with the object union (i.e., contains check for the room and disjoint check for obstacles). If any collision-free pose exists, return True; otherwise False. Compare the model’s Boolean prediction to this result.
Main Issue
The task is harder when non–axis-aligned placements are allowed. Models often mis-handle overlap checks under rotation and falsely report feasibility/infeasibility due to incorrect intersection computations.
G.6 Shortest Path
Task Definition
The question asks: “Determine the shortest valid path that maintains a clearance of $d$ cm from all other objects, starting from centroid of the obj_1 and ending at the centroid of the obj_2.” In Figure 13 (living room), obj_1 is the TV, obj_2 is armchair_2, and $d=15\,\mathrm{cm}$; the goal is to compute the minimum-length collision-free path (and its length). The figure shows a failure case for GPT-OSS-120B, where the predicted path violates the clearance by intersecting the armchair_1.
Ground-Truth Computation
Offset obstacles (equivalently, erode free space) by $0.15\,\mathrm{m}$ to enforce clearance. Run A* on the navigable grid to obtain the shortest collision-free path polyline between TV and armchair_2. A model path is valid if it is collision-free under the same clearance; it is judged correct if its Fréchet distance to the ground-truth path is $\leq 0.6\,\mathrm{m}$.
Main Issue
More objects and overlaps make clearance buffering, merge obstacles, and narrow corridors, increasing failure modes. Models often mishandle overlaps, producing paths that cut through obstacles or declaring no path when one exists.
G.7 Visibility
Task Definition
The prompt is: “Find all objects that intersect the vector from the centroid of the obj_1 to the centroid of the obj_2.” In the visualization in Figure 15 (office), obj_1 is the window and obj_2 is the bin; the goal is to return the set of objects that intersect this segment (excluding the endpoints).
Figure 14: Visibility (office): the table and armchair (highlighted in orange) intersect the line segment from the centroid of the window to the centroid of the bin and constitute the correct answer.
Figure 15: Max Box (bedroom): the red rectangle shows the largest rectangular region that can be placed without overlaps, excluding soft coverings such as rugs.
Ground-Truth Computation
Compute the centroids of window and bin; form the line segment between these centroids. Return the set of objects whose bounding boxes intersect this segment, excluding endpoint touches (i.e., ignore cases where the segment only touches at its endpoints).
Main Issue
On HSSD layouts, accuracy is slightly worse due to the larger number of objects and overlaps; the increased number of polygons along the line increases intersection ambiguity and error rates.
G.8 Max Box
Task Definition
We consider the following question: “Calculate the area in square meters of the largest rectangle that can fit inside the room”. In the visualization in Figure 15 (bedroom), the goal is to compute the maximum-area non–axis-aligned rectangle that fits without overlapping any obstacles, and to report its area.
Ground-Truth Computation
Let $R$ be the room polygon and $O$ the union of all object polygons except rugs (soft coverings). Compute free space $F=R\setminus O$. Search over orientations $\theta\in[0,\pi)$: rotate $F$ by $\theta$, find the largest axis-aligned empty rectangle inside the rotated $F$, record $(w_{\theta},h_{\theta})$ and area $A_{\theta}=w_{\theta}h_{\theta}$, then map back to get $(w^{*},h^{*},\theta^{*})$ with $A^{*}=\max_{\theta}A_{\theta}$.
Main Issue
Harder than simple placement: the model must optimize size and orientation, not just answer yes/no. Allowing rotation makes the search non-convex; more objects and overlaps increase combinatorial complexity. Models often (i) ignore rotation and return an axis-aligned box, or (ii) mis-handle overlaps in free space, leading to under- or over-estimated maxima.
@@ Appendix H Prompts
This section illustrates the design of prompt templates used in our benchmark. We first show a representative example of a question prompt, demonstrating how natural language templates are instantiated to elicit spatial reasoning skills (Figure 16).
Next, we present two examples of layout-generation prompts for bedrooms. The first specifies the creation of base room boundaries and openings (walls and windows) (Figure 17). The second demonstrates how furniture and objects are placed within the generated layout to yield a complete scene (Figure 18).
For completeness, full prompt templates, formatting rules, and implementation details are provided in the supplementary code to support reproducibility.
Prompt: Free Space Given the {room_type} layout in {format}, calculate the total non-occupied (free) floor area in square meters (m2). Room layout: {room} Begin with printing a concise checklist (3–7 bullets) of the conceptual steps necessary for calculating the free space. Then, carefully walk through each reasoning step required to calculate the area. If the format, object names, or required input data are missing, invalid, or inconsistent, reply with: *Final answer*: ERROR Limit your output to the step-by-step reasoning only, and do not include any internal reasoning unless explicitly requested. Clearly state the final answer on the last line using the exact format specified below. ### Output Format
<step-by-step calculations>
*Final answer*: <area>
Where <area> is a float rounded to three decimal places, representing the free area in m2. For example: *Final answer*: 12.347
@@
Prompt: Generate Bedroom Layouts Generate a dataset of {N} bedroom layouts in JSON format. Each layout must include: A unique layout_id A room dictionary with: width, depth, units (meters) shape (”rectangular”, ”L-shaped”, or ”open”) shape_description, intended_use, and bed_size_suggestion An objects list with dictionaries containing: label, bbox [y0, x0, y1, x1], and a descriptive comment Layouts must obey structural and spatial constraints: 50% rectangular, 30% L-shaped, 20% open. L-shape cutouts in corners; each remaining segment $\geq$ 1.5 m. All layouts must include a door (except open types); avoid placing doors and windows on the same short wall. Windows must span >>15% of usable floor area, with equal sizing on shared walls and valid grouping logic. Optional elements: fireplace (for master bedrooms), closet alcove. No overlap or out-of-bound placement. Fireplace must not overlap with doors/windows. Follow a consistent coordinate system: top-left origin, x=width (left to right), y=depth (top to bottom). Return a JSON list of {N} valid layouts. No comments or trailing metadata.
@@
Appendix I Supplementary Accuracy Analyses
Table 10: % token-limit stop reason for general models.
Question Room [Uncaptioned image] claude sonnet-4 [Uncaptioned image] gpt-4.1 [Uncaptioned image] Kimi-K2 Instruct [Uncaptioned image] Qwen3 Coder-480B [Uncaptioned image] Qwen3 235B [Uncaptioned image] gpt-4.1 mini [Uncaptioned image] Qwen3 30B [Uncaptioned image] Devstral Small
Pair distance K 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.2
LR 0.0 0.2 0.0 0.0 0.0 0.0 0.0 0.2
B 0.0 0.0 0.2 0.0 0.0 0.2 0.0 0.8
HSSD 0.0 0.5 0.0 0.0 17.5 8.5 3.5 1.5
Placement K 0.0 0.0 0.0 0.0 6.5 0.0 1.2 1.2
LR 0.0 0.0 0.0 0.0 6.0 0.0 1.0 1.5
B 0.0 0.0 0.0 0.0 17.2 0.0 2.2 0.5
HSSD 0.0 0.0 0.5 0.0 6.5 0.0 1.5 4.5
Reposi- tioning K 0.0 0.2 0.0 0.0 5.7 0.0 2.3 0.2
LR 0.0 0.2 0.0 0.0 1.7 0.0 1.0 1.0
B 0.0 0.0 0.0 0.0 7.5 0.2 0.8 1.3
HSSD 0.0 0.0 0.0 0.0 47.0 2.0 29.0 4.5
Free space K 0.0 0.0 0.2 0.0 0.2 0.0 0.3 2.5
LR 0.0 0.3 0.0 0.0 4.0 0.0 1.0 10.2
B 0.0 0.0 0.0 0.0 4.5 0.0 0.3 6.3
HSSD 0.0 2.0 0.0 0.0 52.5 1.0 39.0 28.5
Visibility K 0.2 0.0 0.0 0.0 0.0 0.0 0.2 0.3
LR 0.2 0.0 0.0 0.0 0.2 0.2 0.5 0.3
B 1.0 0.0 0.0 0.0 0.2 0.3 1.3 0.7
HSSD 0.0 0.0 0.0 0.0 11.5 2.5 11.0 0.0
View angle K 0.0 0.0 0.0 0.0 0.5 0.0 0.0 0.3
LR 0.0 0.0 0.2 0.2 1.5 0.0 0.2 0.7
B 0.0 0.0 0.0 0.3 0.5 0.0 0.5 1.2
HSSD 0.0 0.5 0.0 0.0 26.5 3.0 17.5 2.5
Max box K 0.0 0.0 0.0 0.0 22.5 0.0 6.2 1.5
LR 0.0 0.0 0.7 0.0 49.8 0.0 29.3 0.8
B 0.0 0.0 1.5 0.0 46.5 0.3 21.8 0.8
HSSD 0.0 0.0 0.5 0.0 45.5 0.5 20.0 3.0
Shortest path K 0.0 0.2 0.5 0.0 44.3 0.0 41.5 6.7
LR 0.0 0.0 0.0 0.0 37.8 0.2 46.3 7.8
B 0.2 0.0 0.0 0.0 40.2 0.0 49.0 9.2
HSSD 0.0 0.0 0.0 0.0 35.5 0.0 43.0 27.0
Table 11: Question-level accuracy on completed answers for general models.
Question Room [Uncaptioned image] claude sonnet-4 [Uncaptioned image] gpt-4.1 [Uncaptioned image] Kimi-K2 Instruct [Uncaptioned image] Qwen3 Coder-480B [Uncaptioned image] Qwen3 235B [Uncaptioned image] gpt-4.1 mini [Uncaptioned image] Qwen3 30B [Uncaptioned image] Devstral Small
Pair distance K 99.8 96.5 95.7 96.8 99.2 90.7 89.5 58.9
LR 99.5 95.2 94.2 96.8 99.7 88.5 88.8 62.4
B 99.7 96.3 93.5 96.5 99.5 88.0 85.2 60.5
HSSD 88.0 56.3 75.5 66.0 81.2 40.4 46.1 56.9
Placement K 87.8 78.0 82.2 80.3 96.4 86.5 86.5 75.0
LR 80.5 69.0 73.8 83.2 94.9 75.2 86.5 72.9
B 68.8 59.8 67.5 70.8 92.6 68.7 78.7 56.6
HSSD 72.0 64.5 73.4 70.0 87.7 76.5 72.6 57.1
Reposi- tioning K 73.8 63.9 48.7 45.3 88.5 66.3 75.4 14.9
LR 79.3 60.4 56.7 64.5 92.9 76.8 72.9 25.3
B 71.0 55.5 48.3 59.5 85.4 72.6 70.6 21.5
HSSD 42.0 47.0 28.0 34.0 73.6 41.3 46.5 10.5
Free space K 97.8 93.2 83.1 84.2 95.2 95.2 84.1 67.9
LR 0.2 14.2 2.8 1.8 3.7 1.2 0.5 3.0
B 2.7 31.2 1.0 1.3 9.3 0.8 1.0 0.7
HSSD 35.0 16.3 24.0 22.0 35.8 15.7 12.3 9.1
Visibility K 63.3 87.7 54.5 67.0 98.3 90.2 91.6 18.6
LR 52.8 81.7 43.3 52.2 98.5 87.0 89.1 10.0
B 58.1 74.8 41.0 54.0 96.8 86.6 87.5 10.9
HSSD 20.0 46.5 22.5 26.5 79.7 53.3 51.1 9.0
View angle K 92.0 95.3 69.8 78.8 97.5 95.8 74.7 49.8
LR 87.7 93.2 59.8 75.3 95.4 93.0 72.3 40.8
B 88.0 90.2 60.3 72.8 95.5 91.2 77.2 36.3
HSSD 67.5 55.3 28.5 46.5 68.7 47.4 41.8 30.8
Max box K 47.2 31.8 32.2 26.9 84.5 27.5 28.2 4.6
LR 7.8 7.0 5.1 5.7 44.5 4.5 12.5 0.5
B 5.8 6.8 7.4 5.0 54.5 4.9 9.0 1.7
HSSD 5.0 7.5 2.5 2.0 21.1 4.5 3.8 1.0
Shortest path (valid) K 59.2 61.7 52.7 39.7 81.1 55.3 37.3 36.6
LR 53.0 51.3 42.8 37.3 72.1 47.3 37.9 32.9
B 48.6 52.2 40.7 34.7 73.8 45.8 36.3 36.9
HSSD 28.5 25.0 18.5 18.0 40.3 15.0 9.7 14.4
Shortest path (Fréchet) K 45.3 56.8 51.3 28.2 79.3 39.5 30.5 38.8
LR 24.7 42.5 27.3 12.3 55.8 26.5 17.1 16.8
B 23.0 42.2 27.7 14.2 63.5 30.5 16.0 20.2
HSSD 15.5 22.5 18.0 8.0 31.0 12.5 4.4 15.8
Table 12: % token-limit stop reason for reasoning models.
Question Room [Uncaptioned image] gpt-5 [Uncaptioned image] gpt-oss 120b [Uncaptioned image] DeepSeek R1-0528 [Uncaptioned image] Gemini Flash 2.5 [Uncaptioned image] gpt-5 mini-2025 [Uncaptioned image] gpt-oss 20b [Uncaptioned image] Qwen3 30B Think.
Pair distance K 0.0 0.0 1.5 3.3 0.0 0.7 2.0
LR 0.0 0.0 0.8 4.0 0.2 0.7 2.3
B 0.0 0.0 2.5 4.3 0.2 0.7 4.7
HSSD 0.0 11.0 70.0 86.5 67.0 39.5 72.0
Placement K 13.2 0.0 5.2 39.2 6.5 4.5 29.5
LR 22.7 0.2 7.2 45.8 10.3 12.3 52.5
B 36.2 0.2 8.2 63.3 12.8 25.7 64.5
HSSD 27.0 0.0 6.0 84.0 16.5 15.0 70.5
Reposi- tioning K 0.3 0.0 8.3 1.0 0.0 0.8 7.2
LR 0.0 0.0 7.5 2.2 0.0 0.5 6.7
B 0.2 0.0 3.2 1.7 0.0 0.8 8.3
HSSD 2.0 0.5 33.5 76.0 28.5 22.5 63.0
Free space K 1.2 0.0 5.5 2.0 0.0 0.8 2.3
LR 0.2 0.2 41.7 41.7 1.7 13.7 2.3
B 0.2 0.0 12.2 30.7 1.0 7.8 3.0
HSSD 5.0 28.5 79.5 96.5 85.0 77.5 98.5
Visibility K 0.0 1.5 26.7 72.8 0.0 0.3 19.2
LR 0.0 1.7 46.8 88.2 0.3 1.0 26.3
B 0.0 1.2 44.5 88.3 1.5 0.2 33.3
HSSD 0.0 5.0 87.0 99.5 56.5 29.0 96.0
View angle K 0.0 0.0 25.5 5.5 11.2 1.0 1.0
LR 0.0 0.0 30.8 5.0 14.3 1.3 0.7
B 0.0 0.0 23.8 3.7 12.7 1.2 0.8
HSSD 0.0 7.5 84.5 75.5 73.0 38.0 66.0
Max box K 0.0 0.0 30.2 95.7 2.0 28.8 62.7
LR 0.0 0.0 63.5 100.0 5.7 61.0 98.0
B 0.0 0.0 59.3 100.0 4.7 56.0 98.2
HSSD 0.0 0.0 58.5 100.0 22.5 33.0 99.0
Shortest path K 0.2 0.5 78.5 96.2 23.5 36.2 85.2
LR 61.7 1.2 77.8 97.8 19.8 40.3 81.3
B 0.5 0.5 79.5 98.0 17.5 50.3 83.7
HSSD 0.0 3.0 82.0 100.0 65.5 32.5 98.0
Table 13: Question-level accuracy on completed answers for reasoning models.
Question Room [Uncaptioned image] gpt-5 [Uncaptioned image] gpt-oss 120b [Uncaptioned image] DeepSeek R1-0528 [Uncaptioned image] Gemini Flash 2.5 [Uncaptioned image] gpt-5 mini-2025 [Uncaptioned image] gpt-oss 20b [Uncaptioned image] Qwen3 30B Think.
Pair distance K 99.8 99.3 98.0 96.3 100.0 94.2 97.7
LR 98.8 99.3 99.0 96.0 99.7 93.5 97.5
B 98.3 99.5 96.8 95.5 99.7 93.8 95.3
HSSD 69.0 78.5 25.5 12.5 32.5 40.5 18.0
Placement K 84.7 92.0 89.0 59.7 90.8 85.7 68.2
LR 75.5 89.0 82.2 53.3 86.3 78.5 46.5
B 61.2 83.5 72.0 35.8 81.2 62.0 34.5
HSSD 70.0 85.0 79.0 16.0 75.5 74.5 28.5
Reposi- tioning K 83.0 85.5 79.2 90.5 84.5 70.8 61.5
LR 85.5 89.8 86.2 91.2 92.8 87.8 77.0
B 77.8 83.3 83.3 85.2 84.3 78.0 69.2
HSSD 49.5 60.5 47.5 18.5 53.5 40.0 27.0
Free space K 82.5 99.0 93.0 93.3 99.5 94.8 97.0
LR 47.0 83.3 18.3 17.7 78.5 53.2 0.0
B 50.5 87.5 34.8 33.3 82.2 74.0 1.0
HSSD 19.5 31.0 6.5 1.0 5.0 9.0 1.0
Visibility K 94.8 94.2 71.3 26.8 98.0 91.5 78.8
LR 95.2 94.0 52.0 11.3 98.0 89.3 70.8
B 94.2 92.5 53.5 11.2 95.5 89.2 64.2
HSSD 57.0 70.0 10.0 0.5 39.0 45.5 3.0
View Angle K 96.2 98.5 73.7 92.5 88.3 93.5 98.5
LR 93.3 97.3 68.2 91.5 84.5 92.0 98.3
B 95.2 98.2 75.2 93.8 86.8 91.3 98.3
HSSD 59.5 74.0 13.5 20.0 25.5 37.5 26.0
Max Box K 48.5 62.8 50.5 3.7 85.2 31.3 16.7
LR 17.3 28.2 8.0 0.0 60.3 3.8 0.8
B 13.0 30.3 11.0 0.0 61.2 5.8 0.5
HSSD 5.0 9.5 2.5 0.0 17.0 0.5 0.0
Shortest path (valid) K 64.7 64.2 12.3 3.2 52.3 43.0 14.2
LR 21.2 66.0 12.5 1.5 53.7 37.7 16.7
B 58.2 57.8 7.8 1.0 52.0 30.0 14.3
HSSD 28.5 33.5 8.5 0.0 16.5 13.5 1.0
Shortest path (Fréchet) K 56.3 55.5 11.5 3.2 50.5 42.2 14.0
LR 12.3 40.5 9.3 1.3 47.5 28.5 16.3
B 39.3 44.8 6.0 0.8 47.7 24.2 14.3
HSSD 22.5 26.5 2.5 0.0 12.0 14.0 1.0